OpenStudy (decentnabeel):

A local fish market is selling fish and lobster by the pound. The fish costs $5.00 a pound, while the lobster costs $10.50 a pound. The fish market sells 30 total pounds and makes $194. Represent this situation with a system of equations and solve it to find how many pounds of fish and lobster were sold?

OpenStudy (decentnabeel):

@sweetburger

OpenStudy (decentnabeel):

@ShadowLegendX

OpenStudy (decentnabeel):

x=pounds of fish) y=pounds of lobster 5x+10.50y=194 x+y=30 to solve: No.1 5x+10.50y=194 No.2 x+y=30 subtract y on both sides of No.2 x=−y+30 replace x in No.1 with altered No.2 5(−y+30)+10.50y=194 −5y+150+10.50y=194 5.50y+150=194 5.50y=44 y=8 replace y with 8 in origional No.2 x+8=30 x=22

OpenStudy (xguardians):

I should make a guide on how to solve constraint problems, I see so many of these and I feel like they are so repetitive :P. I believe the work you have done here is correct.